OpenStudy (anonymous):

derivative of (4e)^x

OpenStudy (kainui):

We can simplify this a little bit by using some basic algebra rules of exponents. \[(4e)^x=4^xe^x\] Now you can use the product rule and take the derivative fairly easily!

OpenStudy (anonymous):

\[\LARGE(4e)^xln(4e)=(4e)^xln4+(4e)^x\]

OpenStudy (anonymous):

Hey Cinar is this the answer;

OpenStudy (anonymous):

yes it is..

OpenStudy (anonymous):

Thanks.

Parth (parthkohli):

\[(4e)^x = {4^x}{e^x}\]Applying the product rule, we have\[4^x\ln|4|e^x + 4^xe^x =e^x(4^x\ln|4|+4^x)\]